A small company invested $20,000 by putting part of it into a mutual bond fund that earned 4.5% annual simple interest and the remainder in a corporate bond fund that earned 9.5% annual simple interest. If the company earned $1500 annually from the investments, how much was in the mutual bond fund?
Answer by addingup(3677)

Invested part of 20,000 at 4.5%. Let's call this amount x
The rest of the 20,000 they invested at 9.5%
They earned 1,500/year
--------------------------------
Let's put this information into an equation:
0.045x + 0.095(20,000 - x) = 1,500
0.045x + 1,900 - 0.095x = 1,500
0.045x - 0.095x + 1,900 = 1,500
-0.05x = -400
x = 8,000 So they invested 8,000 at 4.5% and 12,000 at 9.5%
----------------------------------
Check:
8,000(0.045) + 12,000(0.095) = 1,500